OpenStudy (ilovecake):

What is the measure of the missing angle? A. 50° B. 20° C. 40° D. 60°

OpenStudy (k_lynn):

Where is the figure?

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (ilovecake):

there

OpenStudy (alexandervonhumboldt2):

do you know triangle sum theorem?

OpenStudy (alexandervonhumboldt2):

a+b+c=180 70+70+c=180

OpenStudy (alexandervonhumboldt2):

c=180-70-70=????????/

OpenStudy (k_lynn):

you need to know that all the sides of a triangle add up to 180 degrees, as @AlexandervonHumboldt2 said, this is the triangle sum theorem. SO you need to add 70+70 and then subtract that sum by 180.

OpenStudy (k_lynn):